# Longhorn: State Space Models are Amortized Online Learners

### v4

@@ 5.3 Large-scale Language Modeling
For the large-scale language modeling task, we followed the experiment setting in GLA (Yang et al., 2023), where a model with around 1.3B parameters is trained on the SlimPajama (Soboleva et al., 2023) dataset over 100B tokens with a batch size of 2M tokens. We use the AdamW optimizer (Loshchilov & Hutter, 2017) with weight decay of 0.01, and apply cosine learning rate decay with a peak learning rate of $43e-$ 4 and the final learning rate of $53e-$ 5. In addition, we use gradient clipping of 1.0. Due to the large size of the experiment, we only compare against LLaMA, Mamba and GLA models, all with a context size of 2048.333Our result is in general better than those reported in the GLA paper, this might due to the fact that we are using a different 100B subset of SlimPajama since the original split of the SlimPajama is not provided.
We evaluate the trained models on a wide range of downstream tasks covering both common-sense reasoning and question-answering. In particular, we consider PIQA (Bisk et al., 2020), HellaSwag (Hella) (Zellers et al., 2019), WinoGrande (Wino) (Sakaguchi et al., 2021), ARC-easy (ARC-e) and ARC-challenge (ARC-c) (Clark et al., 2018), OpenBookQA (OBQA) (Mihaylov et al., 2018), Social Interaction QA (SIQA) (Sap et al., 2019), and Boolean questions (BoolQ) (Clark et al., 2019).
We report the average perplexity across the above eight datasets throughout training in Figure 1 (left). Then we summarize the downstream evaluation results in Table 4.
@@
LLaMA 1024K 55.08 55.36 71.73 59.26 32.19 43.35 45.16 62.13 53.03
GLA 1024K 55.55 49.10 71.12 58.86 28.11 41.67 44.91 59.21 51.07
Mamba 32K 54.21 53.61 71.67 61.05 30.15 43.94 44.18 59.22 52.25
Longhorn 32K 55.78 52.30 71.00 60.63 29.53 43.55 44.68 61.29 52.35
@@ 5.4 Ablation on Length Extrapolation
Moreover, we evaluate how Longhorn extrapolates to a context length longer than 2048 (training context length) at inference time. In particular, we pick a disjoint validation set from SlimPajama dataset, rearrange it into batches of sequences of length $T\in\{2048,4096,8192,16384,32768\}$, and then evaluate the pretrained model’s perplexity on those sequences. The results are summarized in Figure 1 (right).
Observation: From the figure, we observe that Longhorn successfully extrapolates to contexts up to 16x longer than those used during training, this contrasts with DeltaNet (Yang et al., 2024), which highlights a limitation in that the model cannot extrapolate to longer contexts. Moreover, it is expected that LLaMA, which is a Transformer-based model, fails to extrapolates beyond its training context length.
@@ 6 Conclusion and Future Work
In this work, we introduce a novel approach to designing deep state-space models (SSMs) by conceptualizing the recurrence update as solving an online objective. We propose a straightforward online regression objective, adopting its implicit closed-form update to define our model, which we refer to as Longhorn. Notably, Longhorn is highly parallelizable and achieves state-of-the-art performance in both synthetic sequence modeling and language modeling tasks. For future research, an intriguing avenue would be exploring other online learning objectives that can be efficiently implemented on modern hardware. Additionally, while the current implementation of Longhorn closely aligns with Mamba, recent studies (Ren et al., 2024) suggest that incorporating sliding-window attention with Mamba improves performance. We anticipate similar benefits for Longhorn.

### v5

@@ 5.3 Large-scale Language Modeling
For the large-scale language modeling task, we followed the GLA (Yang et al., 2023) setup, training a 1.3B parameter model on the SlimPajama (Soboleva et al., 2023) dataset with 100B tokens and a batch size of 2M. We used the AdamW optimizer (Loshchilov & Hutter, 2017) with a weight decay of 0.01, cosine learning rate decay (peak: $43e-$ 4, final: $53e-$ 5), and gradient clipping of 1.0. Comparisons were made against LLaMA, Mamba, and GLA models (context size: 2048). We evaluated on eight standard downstream tasks, including PIQA (Bisk et al., 2020), HellaSwag (Hella) (Zellers et al., 2019), WinoGrande (Wino) (Sakaguchi et al., 2021), ARC-easy (ARC-e) and ARC-challenge (ARC-c) (Clark et al., 2018), OpenBookQA (OBQA) (Mihaylov et al., 2018), Social Interaction QA (SIQA) (Sap et al., 2019), and Boolean questions (BoolQ) (Clark et al., 2019). We report the average perplexity across the above eight datasets throughout training in Figure 1 (left). Then we summarize the downstream evaluation results in Table 2.
@@
LLaMA 8M 55.08 55.36 71.73 59.26 32.19 43.35 45.16 62.13 53.03
GLA 512K 55.55 49.10 71.12 58.86 28.11 41.67 44.91 59.21 51.07
Mamba 64K 54.21 53.61 71.67 61.05 30.15 43.94 44.18 59.22 52.25
Longhorn 64K 55.78 52.30 71.00 60.63 29.53 43.55 44.68 61.29 52.35
@@ 5.4 Ablation on Length Extrapolation
We evaluate how Longhorn extrapolates to a context length longer than 2048 (training context length) at inference time. In particular, we pick a disjoint validation set from SlimPajama dataset, rearrange it into batches of sequences of length $T\in\{2048,4096,8192,16384,32768\}$, and then evaluate the pretrained model’s perplexity on those sequences. The results are summarized in Figure 1 (right).
Observation: From the figure, we observe that Longhorn successfully extrapolates to contexts up to 16x longer than those used during training, this contrasts with DeltaNet (Yang et al., 2024), which highlights a limitation in that the model cannot extrapolate to longer contexts. In contrast, LLaMA, as a Transformer-based model, fails to extrapolate beyond its training context length.
5.5 Vision State Space Models
In addition to language tasks, recent works have also applied state space models to the vision domain, leveraging their superior training efficiency. In particular, following the Vision Mamba (ViM) (Zhu et al., 2024), we conduct experiments on the ImageNet (Deng et al., 2009) classification task. Similar to ViM, We apply a bi-directional scan with Longhorn SSM (ViL) and compare the results with ViM on both the tiny and small configurations described in the ViM paper.
Model # Param Top-1 Accuracy
ViM-Tiny 7M 76.1
ViL-Tiny (ours) 7M 76.4
ViM-Small 26M 80.5
ViL-Small (ours) 26M 80.7
Table 3: Top-1 Accuracy on ImageNet classification for Vision Mamba (ViM) and Vision Longhorn (ViL).
Observation: The results from Table 3 demonstrate that the Vision Longhorn model (ViL) achieves comparable (slightly better) performance to the original ViM. Note that we use the best hyperparameters for ViM without additional tuning, and ViL does not require two additional parameters for the forward and backward $A$ matrices, as they are computed directly based on the key $k$ vector.
@@ 6 Conclusion and Future Work
This work introduces a novel approach to designing deep state-space models (SSMs) by conceptualizing the recurrence update as solving an online objective. Based on this, we propose Longhorn, a novel SSM model that explicitly solves online associative recall in closed form. Longhorn is parallelizable and achieves state-of-the-art performance among SSMs on MQAR, language modeling, and image classification tasks. One future direction is to explore other online learning objectives. Additionally, recent studies (Ren et al., 2024) suggest that incorporating sliding-window attention with Mamba improves performance. We anticipate similar benefits for Longhorn.
Appendix A Prior Deep State Space Models
Example A.1 (Linear Attention Variants).
Example A.2 (Mamba (Gu & Dao, 2023)).
$\begin{split}S_{t}&=A_{t}\odot S_{t-1}+B_{t},~{}~{}~{}~{}\text{where}\\ A_{t}&=\exp(A\odot(\varepsilon(x_{t})\otimes 1)),\qquad B_{t}=(\varepsilon(x_{t})\odot x_{t})\otimes k(x_{t}).\end{split}$ (8)
where $S_{t}\in\mathbb{R}^{d\times m}$ with $m=16$ by default, $\varepsilon(x_{t})\in\mathbb{R}^{d}$, $k(x_{t})\in\mathbb{R}^{m}$ linear mappings of $x_{t}$, and $A\in\mathbb{R}^{d\times m}$ is a data independent (not depending on $x_{t}$ trainable weight matrix.
In Mamba, both $A_{t}$ and $B_{t}$ depend on $\varepsilon(x_{t})$, which represents the step size for the SSM update.
Example A.3 (Griffin (De et al., 2024)).
In Mamba and the linear attention variants, the outer product serves as a critical role in lifting vectors to matrices. The recent Griffin architecture abandons the outer product and performs pure elementwise product:
$s_{t}=a(x_{t})\odot s_{t-1}+\sqrt{1-a(x_{t})}\odot i(x_{t})\odot x_{t},$
where $s_{t},a(x_{t}),i(x_{t})$ are all $\mathbb{R}^{d}$. This yields smaller memory states, but in practice, Griffin is combined with local attention (i.e., the sliding-window self-attention) to strengthen its capability.
Example A.4 (RWKV (Peng et al., 2023)).
$s_{t}$ $=u_{t}/z_{t}$
Example A.5 (HGRN2 (Qin et al., 2024a)).
$S_{t}=(1\otimes f(x_{t}))\odot S_{t-1}+i(x_{t})\otimes(1-f(x_{t})).$
Appendix B Existing State Space Models’ Online Objectives
We reverse-engineer some existing deep SSMs’ online learning objectives in Table 4.
LA $\left\lVert S-S_{t-1}\right\rVert^{2}_{\text{F}}-2\langle Sk_{t},x_{t}\rangle$ $S_{t}=S_{t-1}+x_{t}\otimes k_{t}$
RetNet $\gamma\left\lVert S-S_{t-1}\right\rVert^{2}+(1-\gamma)\left\lVert S\right\rVert^{2}_{\text{F}}-2\langle Sk_{t},x_{t}\rangle$ $S_{t}=\gamma S_{t-1}+x_{t}\otimes k_{t}$
GLA $\left\lVert S-S_{t-1}\mathrm{diag}(\alpha_{t})\right\rVert^{2}_{\text{F}}+2\langle Sk_{t},x_{t}\rangle$ $S_{t}=S_{t-1}\mathrm{diag}(\alpha_{t})+x_{t}\otimes k_{t}$
Longhorn $\left\lVert S-S_{t-1}\right\rVert^{2}_{\text{F}}+\left\lVert Sk_{t}-x_{t}\right\rVert^{2}_{\mathrm{diag}(\beta_{t})}$ $S_{t}=(1_{m\times n}-\varepsilon_{t}\otimes k_{t}^{\odot 2})\odot S_{t-1}+$
$(\varepsilon_{t}\odot x_{t})\otimes k_{t},~{}~{}\varepsilon_{t}=\beta_{t}/(1+\beta_{t}k_{t}^{\top}k_{t})$
Appendix C Proof
This section provides the proof for Theorem 6. Given the Longhorn’s objective $S_{t}=\operatorname*{arg\,min}_{S\in\mathbb{R}^{d\times m}}\left\{||S-S_{t-1}||^{2}_{\text{F}}+||Sk_{t}-x_{t}||^{2}_{\mathrm{diag}(\beta_{t})}\right\}$, we have the following theorem:
Theorem C.1.
The closed form solution for $S_{t}$ for objective in Equation 5 is
$S_{t,i}=(I-\varepsilon_{t,i}k_{t}k_{t}^{\top})S_{t-1,i}+\varepsilon_{t,i}k_{t}x_{t,i},~{}~{}~{}\text{where}~{}\varepsilon_{t,i}=\frac{\beta_{t,i}}{1+\beta_{t,i}k_{t}^{\top}k_{t}}\in[0,\infty).$ (9)
Proof.
As the objective in equation 5 is in a quadratic form with respect to $s$, there is a unique minimum. Observe that each row of $S$ (e.g., $S_{i}$) optimizes the objective independently, therefore we can solve the solution row-wise. By setting the derivative of $\nabla_{S_{i}}L_{t}=0$, we have:
$\nabla_{S_{i}}L_{t}=0$ $\Longleftrightarrow(S_{i}-S_{t-1,i})+\beta_{t,i}(S_{i}^{\top}k_{t}-x_{t,i})k_{t}=0$
$\Longleftrightarrow(I+\beta_{t,i}k_{t}k_{t}^{\top})S_{i}=S_{t-1,i}+\beta_{t,i}k_{t}x_{t,i}$
$\underbrace{\Longleftrightarrow}_{(3)}S_{i}=\bigg{(}I-\frac{\beta_{t,i}}{1+\beta_{t,i}k_{t}^{\top}k_{t}}k_{t}k_{t}^{\top}\bigg{)}S_{t-1,i}+\bigg{(}I-\frac{\beta_{t,i}}{I+\beta_{t,i}k_{t}^{\top}k_{t}}k_{t}k_{t}^{\top}\bigg{)}\beta_{t,i}k_{t}x_{t,i}$
$\Longleftrightarrow\bigg{(}I-\frac{\beta_{t,i}}{I+\beta_{t,i}k_{t}^{\top}k_{t}}k_{t}k_{t}^{\top}\bigg{)}S_{t-1,i}+\frac{(I+\beta_{t,i}k_{t}^{\top}k_{t}-\beta_{t,i}k_{t}k_{t}^{\top})\beta_{t,i}k_{t}x_{t,i}}{I+\beta_{t,i}k_{t}^{\top}k_{t}}$
$\underbrace{\Longleftrightarrow}_{(5)}\bigg{(}I-\frac{\beta_{t,i}}{I+\beta_{t,i}k_{t}^{\top}k_{t}}k_{t}k_{t}^{\top}\bigg{)}S_{t-1,i}+\frac{\beta_{t,i}k_{t}x_{t,i}}{I+\beta_{t,i}k_{t}^{\top}k_{t}}$
(3) is derived from the fact that $(I+\beta_{t,i}k_{t}k_{t}^{\top})^{-1}=(I-\frac{\beta_{t,i}k_{t}k_{t}^{\top}}{1+\beta_{t,i}k_{t}^{\top}k_{t}})$ by the Sherman–Morrison formula. (5) is derived by noticing that $k_{t}^{\top}k_{t}k_{t}x_{t,i}-k_{t}k_{t}^{\top}k_{t}x_{t,i}=0$. ∎
Appendix D Additional Experiment Details
We provide the architecture detail for conducting the scaling law experiments on OpenWebText in Table 5. The architecture configs follow exactly from the Mamba paper (Gu & Dao, 2023).
125M 12 768 12 / 64 4800 6e-4 0.5M tokens 2.5B
350M 24 1024 16 / 64 13500 3e-4 0.5M tokens 7B
Table 5: Training details on OpenWebText.